Determine the test criteria:

Specify the level of significance (Type I error associated with the null hypothesis),

Determine the test statistic (the appropriate statistical test as mentioned under point 5 above),

Determine the critical values (and region(s) if applicable),

Calculate the value of the test statistic (obtained value).

Make a decision about the null and research hypothesis by comparing the obtained value to the critical value and interpret the results of the data. You can pay attention also to the p-values.

ANSWER IN EXCEL FORM

Shift Stress
1 2
1 5
1 7
1 8
1 5
1 4
1 3
1 6
1 5
1 4
2 4
2 5
2 3
2 6
2 16
2 11
2 9
2 8
2 12
3 7
3 13
3 10
3 16
3 11
3 11
3 7
3 3
3 10
3 15
3 12
Answer #1

Solution

The solution is obtained through ANOVA (Analysis of Variance) – One-way with three shifts as factor.

Let xij be the jth stress observation in the ith shift, i = 1, 2, 3 and j = 1 to ni, n1 = 10, n2 = 9, n3 = 11.

Then the ANOVA model is: xij = µ + αi + εij, where µ = common effect, αi = effect of ith shift, and εij is the

error component which is assumed to be Normally Distributed with mean 0 and variance σ2.

Null hypothesis: H0: α1 = α2 = α3 Vs Alternative: HA: H0 is false [at least one αi is different from others]

Final ANOVA Table and decisions and conclusions are given below. Detailed Back-up Theory and Excel

calculations follow at the end.

ANOVA

Table

alpha

0.05

Source

DF

SS

MS

Fobs

Fcrit

p-value

Shift

2

162.68

81.34192

7.012619

3.354131

0.0035255

Error

27

313.18

11.59936

Total

29

475.87

16.4092

Decision: Since Fobs > Fcrit or equivalently, p-value < alpha, H0 is rejected.

Conclusion: There is enough evidence to suggest that the mean stress levels are different for the 3 shifts. Answer

Now, to work out the solution,

Back-up Theory

Terminology:

Shift total = xi.= sum over j of xij

Grand total = G = sum over i of xi.

Correction Factor = C = G2/N, where N = total number of observations = n1 + n2 + n3

Total Sum of Squares: SST = (sum over i,j of xij2) – C

Shift Sum of Squares: SSR = {(sum over i of xi.2)/(ni)} – C

Error Sum of Squares: SSE = SST – SSR

Mean Sum of Squares = Sum of squares/Degrees of Freedom

Fcrit: upper α% point of F-Distribution with degrees of freedom υ1 and υ2, where υ1 is the DF for shift

and υ1 is the DF for Error

Significance: Fobs is significant if Fobs > Fcrit

Calculations

Data

j

i

Sh1

Sh2

Sh3

1

2

4

7

2

5

5

13

3

7

3

10

4

8

6

16

5

5

16

11

6

4

11

11

7

3

9

7

8

6

8

3

9

5

12

10

10

4

-

15

11

-

-

12

Excel Calculations

alpha

0.05

#treat

3

n1

10

n2

9

n3

11

n

30

x1.

49

x2.

74

x3.

115

G = x..

238

C

1888.13

Sx1j^2

269

Sx2j^2

752

Sx3j^2

1343

Sxij^2

2364

Sxi.^2/ni

2050.82

SST

475.867

SSR

162.684

SSE

313.183

DONE